Tom invested 13,000.00 for one year, part at 9 percent annual interest and the rest at 7 percent annual interest. How much must he invest at 9 percent to get $1000.00 in interest in a year?
Found 3 solutions by sujatha.krishna, Tanzina ahamed, Alan3354:
Answer by sujatha.krishna(11)

Interest = Principal *%Rate * Number of years
So 1000 = x * 9/100 * 1
x * 9 * 1
----------- = 1000
100
x = 100000/9
x= 11111.11
So he must invest 11,111.11 for a year at 9% rate to get 1000.00 as interest !!

Answer by Tanzina ahamed(5)

Sujata u have answered wrong in this soln.If anyone want the steps pls contact with me.Anyways the right answer is 4500.

Answer by Alan3354(69443)

Tom invested 13,000.00 for one year, part at 9 percent annual interest and the rest at 7 percent annual interest. How much must he invest at 9 percent to get $1000.00 in interest in a year?
----------
Amt = amount at 9%
Amt*0.09 + (13000 - Amt)*0.07 = 1000
0.02*Amt + 910 = 1000
Amt = 90/0.02 = $4500